Consider a population of 50 individuals with a gene with two alleles, C & c... where "C" is dominant to "c". If two individuals have the "c" phenotype, what is the allele frequency for "C"?

Group of answer choices (This question allows me to answer for more than one answer choice but I'm not sure if there is really more than one)

A) 0.2

B) 0.48

C) 0.8

D) 0.96
Answer #1

Phenotype 1 = CC, Cc

Phenotype 2 = cc = 2 individuals

Genotypic frequency of cc = 2/50 = 0.04

Allelic frequency of c = √0.04 = 0.2

Allelic frequency of C = 1 - 0.2 = 0.8

Option C

Please rate high.